OpenStudy (anonymous):

find the value of y for the given value of x y=x+1, x=8

OpenStudy (imtiaz7):

find the value of y for the given value of x y=x+1 eq. i x=8 eq. ii so simple: put the value of eq ii in eq. i we get: y=8+1 y=9

HanAkoSolo (jamierox4ev3r):

if x=8 then y=(8)+1 therefore, y=9. this is a more simple way of explaining it @uniquejamaica . Does this make sense to you?